Find the standard deviation o for the binomial distribution which has the stated value of n and p round to nearest hundredth n=45 p= 0.2
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

Standard Deviation

sigma = sqrt(np(1-p))

sigma = sqrt(45*0.2(1-0.2))

sigma = 2.68328

sigma = 2.68

So the standard deviation is roughly 2.68

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by stanbon(75887)

Find the standard deviation o for the binomial distribution which has the stated value of n and p round to nearest hundredth n=45 p= 0.2
----
q = 0.8
----
sigma = sqrt(npq) = sqrt(45*0.2*0.8) = 2.6833
==========
Cheers,
Stan H.
===============